# HET/HPF observations of Helium in warm, hot, and ultra-hot Jupiters

### v1

@@ 4.2 Helium circumstellar structures and stellar variability
Following Z. Zhang et al. (2023) and M. Gully-Santiago et al. (2024), we computed the He i triplet line stellar light curve to inspect the presence of circumstellar structures. Figures 13 and 14 do not show evidence of extremely extended outflows of planetary origin, like in HAT-P-32 b and HAT-P-67 b. If such outflows existed, the tails would extend for only a small fraction of the total planetary orbit and/or would not be well sampled. Alternatively, the signal could be diluted within the stellar variability.
@@
We find that stellar helium variability is typically below or of the order of 1 %, ranging from 0.5 % to 2–4 % in some extreme cases. These results are consistent with previous work by D. M. Krolikowski et al. (2024) on He variability with HPF. They monitored the stellar He i triplet line for a sample of young host stars and found that stars older than 300 Myr seem to behave similarly to mature and old stars. These young stars exhibited intrinsic stellar variability of 1 %, ranging 0–2 %. The HET/HPF observational methodology can be used to study atmospheres of exoplanets, even those around moderately young stars.
4.3 Constraints on mass loss rates and helium absorption signals
@@
This work presents the first analyses of the He i triplet for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Although their transmission spectra do not show evidence of planetary absorption, we can estimate the expected helium absorption signal based on the empirical relations from J. Orell-Miquel et al. (2024, see Sect. 6.3). However, we need to compute or estimate the X-ray (5–100 Å) and extreme ultraviolet (EUV, 100–504 Å) irradiation that the planets receive from their host star. Since there are no observations of these new targets in the XUV wavelengths444In this section, XUV refers to wavelengths of X-rays and EUV defined as between 5–504 Å., we compared their stellar properties with those of the available stars in the MUSCLES survey (K. France et al., 2016) and MEATS program (P. R. Behr et al., 2023) to find a suitable comparison star. Then, we scaled the fluxes from the proxy star to our target, accounting for their stellar radii and distances. All MUSCLES and MEATS products are available on the Barbara A. Mikulski Archive for Space Telescopes555https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html (MAST). We want to emphasize that since we scale the target’s luminosity from a proxy star, these results should be considered an order-of-magnitude approximation.
@@
HAT-P-12 ( = 0.701 R⊙, J. D. Hartman et al., 2009) is part of the MEATS survey and its analysis is presented in P. R. Behr et al. (2023). However, the star was not observed at wavelengths below 2500 Å due to its distance ( = 143 pc, C. Soubiran et al., 2018) and low activity. P. R. Behr et al. (2023, Table 2) reports a 3 upper limit of 8.791027 erg s-1 for the X-ray luminosity . We derive 2.401028 erg s-1 using the luminosity relation between X-rays and EUV from J. Sanz-Forcada et al. (2025). Those upper limits are consistent with the luminosities from the HAT-P-12 MUSCLE model, where the X-ray and far ultraviolet fluxes are based on MUSCLES observations of HD85512. We computed 7.121026 erg s-1, and 7.811027 erg s-1. Using the from MUSCLES and the XUV and He i triplet signal relation from J. Orell-Miquel et al. (2024), we obtain an EW = 223 mÅ, consistent with our derived upper limit (21.2 mÅ). The noise level reached by our observations is below this limit. Therefore, an uncontaminated transit of HAT-P-12 b could accurately measure its mass loss. If we consider the upper limit XUV values, we estimate a much larger signal (EW 36 mÅ) which is not supported by our observations.
@@
HAT-P-17 ( = 0.838 R⊙, A. W. Howard et al., 2012; = 92.4 pc, C. Soubiran et al., 2018) is not part of the MUSCLES survey, so we estimated its XUV flux scaling the luminosities of HD97658 ( = 0.728 R⊙, T. G. Ellis et al., 2021; = 21.1 pc, C. Koen et al., 2010) and HAT-P-26 ( = 0.79 R⊙, = 141.8 pc, J. D. Hartman et al., 2011). We computed a 2.51028 erg s-1 from the scaled HD97658 MUSCLES spectrum (R. O. P. Loyd et al., 2016), while 2.51027 erg s-1 is obtained when using the HAT-P-26 MEATS spectrum (P. R. Behr et al., 2023). According to these , we computed an expected He i triplet signal of 3–10 mÅ, which is consistent with the observational upper limit of 7.4 mÅ.
@@
For HD118203 ( = 2.03 R⊙, G. Maciejewski et al., 2024; = 92.04 pc, C. Soubiran et al., 2018), we used as a proxy HD149026 ( = 1.34 R⊙, = 75.0 pc, B. Sato et al., 2005). We computed a 1.261028 erg s-1 from the scaled HD149026 MEATS spectrum (P. R. Behr et al., 2023). Although our observations imposed a restrictive upper limit on the presence of He i (EW 2.1 mÅ), the expected absorption is estimated to be of the same order (2 mÅ).
@@
For TrES-1 ( = 0.807 R⊙, A. S. Bonomo et al., 2017; = 159.622 pc, C. Soubiran et al., 2018), we estimated its XUV flux using HD97658 as a proxy. We computed a 2.351028 erg s-1 from the scaled HD97658 MUSCLES spectrum. From the XUV flux and the empirical relation from J. Orell-Miquel et al. (2024), we derive an EW 91.5 mÅ. The expected EW is consistent and a bit below the upper limit imposed from the observations (11.1 mÅ).
@@
For WASP-156 b ( = 0.826 R⊙, A. S. Polanski et al., 2024; = 121.843 pc, C. Soubiran et al., 2018), we scaled the HAT-P-26 MEATS spectrum as a proxy. We obtained a 2.461027 erg s-1. We estimated an EW absorption of 31 mÅ, well bellow the upper limit from our observations (9.5 Å). Therefore, more transit observations are required to determine more restrictive limits to WASP-156 b’s mass loss.
4.4 Helium detections and trends in the gas giant population
@@
We combined the results on gas giants from this work, from J. Orell-Miquel et al. (2024), and from the ExoAtmospheres database666https://research.iac.es/proyecto/exoatmospheres/index.php; Accessed on 1 August 2025.. Our goal is to explore whether the He i triplet detections in the gas giant population are related to planetary and stellar properties. We obtained a list of 46 planets with 0.5 RJup and He i studies. Of these, less than one-third (14) showed evidence of He absorption.
@@
Figure 7 presents the planet radius of gas giants with helium studies as a function of their host star’s effective temperature (), , and mass. There are no He i detections for gas giants orbiting stars with higher than 6750 K. Half of the detections are in systems with 5000 K, whose planets have 1000 K. The four planets (HAT-P-32 b, Z. Zhang et al., 2023; HAT-P-67 b, M. Gully-Santiago et al., 2024; WASP-76 b, and WASP-121 b, S. Czesla et al., 2024) showing He i outflow orbiting 6750 K stars are also among the largest sampled planets ( 1.6 RJup). These four planets have of 2000 K, but below the UHJ limit (2150 K). No helium detections have been made for exoplanets more massive than 1 MJup.
@@
Figure 7 depicts the three different scenarios for warm, hot, and ultra-hot Jupiter-sized planets. The of the host star, used as a proxy for the spectral type, plays a major role in the detectability of helium. The detection trend for WJs may be similar to that of the general exoplanet population (J. Orell-Miquel et al., 2024). WJs mainly orbit 5000 K host stars, e.g. K-type stars. These stars have a spectral energy distribution that favors the population of the He i metastable state and its detection (A. Oklopčić, 2019). UHJs and some HJs typically orbit hotter stars than WJs, making them less suitable for populating metastable helium. Therefore, their non-detections are likely related to the stellar type of their host rather than a lack of helium atoms in their atmosphere. Inspection of evaporation through other spectral lines, such as H, would be more appropriate for UHJs orbiting A-type stars J. Orell-Miquel et al. (2024). However, the extreme irradiation of F-type stars appears to be enough to make the He i outflow detectable for low-density HJ planets (0.5 g cm-3), such as HAT-P-32 b, HAT-P-67 b, WASP-76 b, and WASP-121 b. Low-density, irradiated gas giants are expected to evaporate at a low rate until they reach a bulk density of 0.1 g cm-3. At this point, they undergo runaway mass loss, which strips away their outer atmospheric layers (D. P. Thorngren et al., 2023). This large outflow would be detectable even if the metastable population mechanism is not efficient, such as in the F-type stars environment. This evaporation process is more intense during the first few hundred Myr, but it can act at Gyr time scales as well (see animation of D. P. Thorngren et al., 2023, Fig 3). All four HJs with helium detections orbit stars older than 1 Gyr.
@@
We also explored the stellar age parameter, but we have not found a clear relation with the He i triplet detections. There are only three gas giants with helium detection around stars younger than 1 Gyr (TOI-5398 b, M. C. D’Arpa et al., 2024; TOI-1268 b, J. Orell-Miquel et al., 2024; and WASP-52 b), but there are also less studied planets, 12 in total. Thus, the ratio of detections to non-detections in the young planet population is similar to that of the old population.
5 Conclusions
@@
For the first time, we provide constraints on the helium line of HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. For these targets, our observational upper limits are consistent with the empirical helium predictions from estimated XUV fluxes. Additionally, we do not detect significant absorption in HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b, which is consistent with previous studies. HAT-P-26 b had a previous detection using narrow-band photometric observations (S. Vissapragada et al., 2022) that we are unable to reproduce with our high-resolution observations.
@@
We compiled a list of 46 irradiated gas giants for which helium studies were conducted, combining the results presented here with those available in the literature. Using this dataset, we explore He i triplet trends within the populations of warm, hot, and ultra-hot Jupiters. We note differences in the detectability of helium among these three populations. Warm Jupiter detections seem to be more closely linked to the host star type, with all of the detections found in planets orbiting K-type stars, which are considered the most favorable stellar type according to theory (A. Oklopčić, 2019). Although there are only four hot Jupiter detections, they come from low-density planets orbiting F-type stars. To date, no helium absorption signal has been reported for ultra-hot Jupiters. Unfortunately, our results are based in limited number of explored atmospheres and further observations of irradiated gas giants are needed to increase this sample and improve our understanding of atmospheric mass loss through He i triplet observations.
Appendix A Planetary system parameters

### v2

@@ IV.2 Helium circumstellar structures and stellar variability
Following Zhang et al. (2023) and Gully-Santiago et al. (2024), we computed the He i triplet line stellar light curve to inspect the presence of circumstellar structures (e.g., Zhang et al., 2023; Gully-Santiago et al., 2024; Krishnamurthy et al., 2025; Allart et al., 2025b). Figures 9 and 10 do not show evidence of extremely extended outflows of planetary origin, like in HAT-P-32 b and HAT-P-67 b. If such outflows existed, the tails would extend for only a small fraction of the total planetary orbit and/or would not be well sampled. Alternatively, the signal could be diluted within the stellar variability.
@@
We find that stellar helium variability is typically below or of the order of 1 %, ranging from 0.5 % to 2–4 % in some extreme cases. These results are consistent with previous work by Krolikowski et al. (2024) on He variability with HPF. They monitored the stellar He i triplet line for a sample of young host stars and found that stars older than 300 Myr seem to behave similarly to mature and old stars. These young stars exhibited intrinsic stellar variability of 1 %, ranging 0–2 %. The HET/HPF observational methodology can be used to study atmospheres of exoplanets, even those around moderately young stars.
IV.3 Constraints on mass loss rates and helium absorption signals
@@
This work presents the first analyses of the He i triplet for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Although their transmission spectra do not show evidence of planetary absorption, we can estimate the expected helium absorption signal based on the empirical relations from Orell-Miquel et al. (2024, see Sect. 6.3). However, we need to compute or estimate the X-ray (5–100 Å) and extreme ultraviolet (EUV, 100–504 Å) irradiation that the planets receive from their host star. Since there are no observations of these new targets in the XUV wavelengths555In this section, XUV refers to wavelengths of X-rays and EUV defined as between 5–504 Å., we compared their stellar properties with those of the available stars in the MUSCLES survey (France et al., 2016b) and its extension for JWST atmospheric transmission spectroscopy targets (Behr et al., 2023) to find a suitable comparison star. Then, we scaled the fluxes from the proxy star to our target, accounting for their stellar radii and distances. All MUSCLES data products are available on the Barbara A. Mikulski Archive for Space Telescopes (MAST) at France et al. (2016a)666https://doi.org/10.17909/T9DG6F (catalog 10.17909/T9DG6F). We want to emphasize that since we scale the target’s luminosity from a proxy star, these results should be considered an order-of-magnitude approximation.
@@
HAT-P-12 ( = 0.701 R⊙, Hartman et al., 2009) is part of the MUSCLES extension survey and its analysis is presented in Behr et al. (2023). However, the star was not observed at wavelengths below 2500 Å due to its distance ( = 143 pc, Soubiran et al., 2018) and low activity. Behr et al. (2023, Table 2) reports a 3 upper limit of 8.791027 erg s-1 for the X-ray luminosity . We derive 2.401028 erg s-1 using the luminosity relation between X-rays and EUV from Sanz-Forcada et al. (2025). Those upper limits are consistent with the luminosities from the HAT-P-12 MUSCLE model, where the X-ray and far ultraviolet fluxes are based on MUSCLES observations of HD85512. We computed 7.121026 erg s-1, and 7.811027 erg s-1. Using the from MUSCLES and the XUV and He i triplet signal relation from Orell-Miquel et al. (2024), we obtain an EW = 223 mÅ, consistent with our derived upper limit (21.2 mÅ). The noise level reached by our observations is below this limit. Therefore, an uncontaminated transit of HAT-P-12 b could accurately measure its mass loss. If we consider the upper limit XUV values, we estimate a much larger signal (EW 36 mÅ) which is not supported by our observations.
@@
HAT-P-17 ( = 0.838 R⊙, Howard et al., 2012; = 92.4 pc, Soubiran et al., 2018) is not part of the MUSCLES survey, so we estimated its XUV flux scaling the luminosities of HD97658 ( = 0.728 R⊙, Ellis et al., 2021; = 21.1 pc, Koen et al., 2010) and HAT-P-26 ( = 0.79 R⊙, = 141.8 pc, Hartman et al., 2011). We computed a 2.51028 erg s-1 from the scaled HD97658 MUSCLES spectrum (Loyd et al., 2016), while 2.51027 erg s-1 is obtained when using the HAT-P-26 spectrum (Behr et al., 2023). According to these , we computed an expected He i triplet signal of 3–10 mÅ, which is consistent with the observational upper limit of 11.9 mÅ.
@@
For HD118203 ( = 2.03 R⊙, Maciejewski et al., 2024; = 92.04 pc, Soubiran et al., 2018), we used as a proxy HD149026 ( = 1.34 R⊙, = 75.0 pc, Sato et al., 2005). We computed a 1.261028 erg s-1 from the scaled HD149026 spectrum (Behr et al., 2023). Although our observations imposed a restrictive upper limit on the presence of He i (EW 2.1 mÅ), the expected absorption is estimated to be of the same order (2 mÅ).
@@
For TrES-1 ( = 0.807 R⊙, Bonomo et al., 2017; = 159.622 pc, Soubiran et al., 2018), we estimated its XUV flux using HD97658 as a proxy. We computed a 2.351028 erg s-1 from the scaled HD97658 MUSCLES spectrum. From the XUV flux and the empirical relation from Orell-Miquel et al. (2024), we derive an EW 91.5 mÅ. The expected EW is consistent and a bit below the upper limit imposed from the observations (13.8 mÅ).
@@
For WASP-156 b ( = 0.826 R⊙, Polanski et al., 2024; = 121.843 pc, Soubiran et al., 2018), we scaled the HAT-P-26 spectrum as a proxy. We obtained a 2.461027 erg s-1. We estimated an EW absorption of 31 mÅ, well bellow the upper limit from our observations (9.5 Å). Therefore, more transit observations are required to determine more restrictive limits to WASP-156 b’s mass loss.
IV.4 Helium detections and trends in the gas giant population
@@
We combined the results on gas giants from this work, from Orell-Miquel et al. (2024), and from the ExoAtmospheres database777https://research.iac.es/proyecto/exoatmospheres/index.php; Accessed on 1 August 2025.. Our goal is to explore whether the He i triplet detections in the gas giant population are related to planetary and stellar properties. We obtained a list of 46 planets with 0.5 RJup and He i studies. Of these, less than one-third (14) showed evidence of He absorption.
@@
Figure 7 presents the planet radius of gas giants with helium studies as a function of their host star’s effective temperature (), , and mass. There are no He i detections for gas giants orbiting stars with higher than 6750 K. Half of the detections are in systems with 5000 K, whose planets have 1000 K. The four planets (HAT-P-32 b, Zhang et al., 2023; HAT-P-67 b, Gully-Santiago et al., 2024; WASP-76 b, and WASP-121 b, Czesla et al., 2024) showing He i outflow orbiting 6750 K stars are also among the largest sampled planets ( 1.6 RJup). HAT-P-32 b and HAT-P-67 b have slightly below the UHJ limit of 2000 K, while WASP-76 b, and WASP-121 b can be considered within the UHJ category. No helium detections have been made for exoplanets more massive than 1.1 MJup.
@@
Figure 7 depicts the three different scenarios for warm, hot, and ultra-hot Jupiter-sized planets. The of the host star, used as a proxy for the spectral type and the spectral energy distribution in the X-ray and UV range, plays a major role in the detectability of helium. The detection trend for WJs may be similar to that of the general exoplanet population (Orell-Miquel et al., 2024). WJs mainly orbit 5000 K host stars, e.g. K-type stars. These stars have a spectral energy distribution that favors the population of the He i metastable state and its detection (Oklopčić, 2019). UHJs and some HJs typically orbit hotter stars than WJs, making them less suitable for populating metastable helium. Therefore, their non-detections are likely related to the stellar type of their host rather than a lack of helium atoms in their atmosphere. Inspection of evaporation through other spectral lines, such as H, would be more appropriate for UHJs orbiting A-type stars (Orell-Miquel et al., 2024). However, the balance between XUV and near-UV irradiation for F-type stars appears to be enough to make the He i outflow detectable for low-density HJ and UHJ planets (0.5 g cm-3), such as HAT-P-32 b, HAT-P-67 b, WASP-76 b, and WASP-121 b. These low-density, irradiated gas giants are expected to evaporate at a low rate until they reach a bulk density of 0.1 g cm-3. At that point, they undergo runaway mass loss, which strips away their outer atmospheric layers (Thorngren et al., 2023). This large outflow would be detectable even if the metastable population mechanism is not efficient, such as in the F-type stars environment (Oklopčić, 2019). This evaporation process is more intense during the first few hundred Myr (e.g., Sanz-Forcada et al., 2011; Owen and Wu, 2013; Gupta and Schlichting, 2020), but it can act at Gyr time scales as well (e.g., see animation of Thorngren et al., 2023, Fig 3). All four HJs and UHJs with helium detections orbit stars older than 1 Gyr.
@@
We also explored the stellar age parameter, but we have not found a clear relation with the He i triplet detections. There are only three gas giants with helium detection around stars younger than 1 Gyr (TOI-5398 b, D’Arpa et al., 2024; TOI-1268 b, Orell-Miquel et al., 2024; and WASP-52 b), but there are also less studied planets, 12 in total. Thus, the ratio of detections to non-detections in the young planet population is similar to that of the old population.
V Conclusions
@@
For the first time, we provide constraints on the helium line of HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. For these targets, our observational upper limits are consistent with the empirical helium predictions from estimated XUV fluxes. Additionally, we do not detect significant absorption in HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b, which is consistent with previous studies. HAT-P-26 b had a previous detection using narrow-band photometric observations (Vissapragada et al., 2022) that we are unable to reproduce with our high-resolution observations.
@@
We compiled a list of 46 irradiated gas giants for which helium studies were conducted, combining the results presented here with those available in the literature. Using this dataset, we explore He i triplet trends within the populations of warm, hot, and ultra-hot Jupiters. We note differences in the detectability of helium among these three populations. Warm Jupiter detections seem to be more closely linked to the host star type, with all of the detections found in planets orbiting K-type stars, which are considered the most favorable stellar type according to theory (Oklopčić, 2019). Although there are only four hot Jupiter detections, they come from low-density planets orbiting F-type stars. To date, no helium absorption signal has been reported for ultra-hot Jupiters. Unfortunately, our results are based in limited number of explored atmospheres and further observations of irradiated gas giants are needed to increase this sample and improve our understanding of atmospheric mass loss through He i triplet observations.
Appendix A Extra tables
| Planet | Date | Class | Num | Phase | Exp. Time | S/N |
| HAT-P-3 b | 2021- 1- 2 | P4 | 4 | 0.32 | 618 | 109 |
| 2021- 1-31 | OUT1 | 4 | 0.35 | 618 | 94 | |
| 2021- 2- 1 | IN1 | 8 | 0.01 | 639 | 95 | |
| 2021- 2- 2 | OUT1 | 4 | 0.34 | 618 | 87 | |
| HAT-P-12 b | 2019-12-18 | P4 | 2 | 0.35 | 916 | 96 |
| 2020- 3-26 | OUT1 | 2 | 0.38 | 777 | 34 | |
| 2020- 3-30 | OUT1 | 1 | 0.32 | 958 | 31 | |
| 2020- 3-31 | OUT1 | 3 | 0.07 | 958 | 32 | |
| 2020- 3-31 | IN1 | 5 | 0.0 | 958 | 57 | |
| 2020- 4- 1 | OUT1 | 4 | 0.31 | 958 | 54 | |
| 2025- 5-26 | OUT2 | 4 | 0.32 | 618 | 60 | |
| 2025- 5-27 | IN2 | 8 | 0.0 | 618 | 47 | |
| HAT-P-17 b | 2020- 5- 7 | P4 | 4 | 309 | 86 | |
| 2020- 6- 7 | OUT1 | 4 | 309 | 77 | ||
| 2020- 6- 8 | IN1 | 11 | 309 | 73 | ||
| 2020- 6-10 | OUT1 | 4 | 309 | 88 | ||
| 2020- 6-11 | OUT1 | 4 | 309 | 76 | ||
| 2020- 7-23 | P4 | 2 | 511 | 102 | ||
| 2020- 7-23 | P4 | 2 | 511 | 159 | ||
| 2020- 7-28 | OUT2 | 4 | 309 | 122 | ||
| 2020- 7-29 | OUT2 | 4 | 309 | 122 | ||
| 2020- 7-30 | IN2 | 10 | 309 | 109 | ||
| 2020- 7-30 | OUT2 | 4 | 309 | 127 | ||
| 2020- 7-31 | OUT2 | 4 | 309 | 105 | ||
| 2020- 8- 1 | OUT2 | 4 | 309 | 99 | ||
| 2020- 8- 1 | OUT2 | 2 | 511 | 141 | ||
| HAT-P-26 b | 2021- 1-31 | OUT1 | 4 | 660 | 96 | |
| 2021- 2- 1 | IN1 | 5 | 660 | 87 | ||
| 2021- 2- 2 | OUT1 | 4 | 660 | 83 | ||
| 2021- 2-24 | P4 | 2 | 820 | 110 | ||
| HD118203 b | 2021- 1- 5 | OUT1 | 2 | 309 | 304 | |
| 2021- 1- 6 | OUT1 | 2 | 309 | 215 | ||
| 2021- 1- 7 | IN1 | 18 | 309 | 344 | ||
| 2021- 1- 8 | OUT1 | 2 | 309 | 331 | ||
| 2021- 1- 9 | OUT1 | 2 | 309 | 290 | ||
| 2021- 1-26 | P4 | 2 | 213 | 228 | ||
| 2021- 2-19 | IN2 | 18 | 309 | 305 | ||
| 2021- 2-19 | OUT2 | 9 | 309 | 322 | ||
| 2021- 2-20 | OUT2 | 2 | 309 | 328 | ||
| HD189733 b | 2019- 4-17 | OUT4 | 11 | 309 | 395 | |
| 2019- 5- 4 | IN1 | 9 | 309 | 403 | ||
| 2019- 5-12 | P4 (OUT1) | 2 | 213 | 320 | ||
| 2019- 6-13 | IN2 | 9 | 309 | 292 | ||
| 2019-10-30 | P4 | 2 | 618 | 737 | ||
| 2020- 7-24 | P4 | 2 | 213 | 402 | ||
| 2020- 8- 1 | IN3 | 2 | 106 | 231 | ||
| 2020- 8-19 | P4 | 2 | 213 | 479 | ||
| 2020-11-24 | P4 | 2 | 213 | 377 | ||
| HD209458 b | 2019-10- 1 | OUT1 | 1 | 447 | 315 | |
| 2019-10- 2 | IN1 | 5 | 522 | 478 | ||
| 2019-10- 4 | OUT2 | 2 | 522 | 159 | ||
| 2019-10- 5 | OUT2 | 4 | 309 | 233 | ||
| TrES-1 b | 2019- 8-12 | P4 | 2 | 618 | 110 | |
| 2020- 7- 5 | OUT1 | 1 | 799 | 96 | ||
| 2020- 7- 7 | IN1 | 5 | 799 | 104 | ||
| 2020- 7- 8 | OUT1 | 4 | 799 | 63 | ||
| 2020- 7-10 | OUT1&2 | 1 | 820 | 93 | ||
| 2020- 7-11 | OUT2 | 4 | 820 | 81 | ||
| 2020- 7-12 | OUT2 | 4 | 820 | 97 | ||
| 2020- 7-13 | OUT2 | 5 | 820 | 99 | ||
| 2020- 7-13 | IN2 | 5 | 820 | 93 | ||
| 2020- 7-14 | OUT2 | 4 | 820 | 88 | ||
| 2020- 7-23 | P4 | 2 | 820 | 116 | ||
| 2020- 8- 1 | P4 | 1 | 820 | 85 | ||
| 2020-10-13 | P4 | 2 | 820 | 91 | ||
| 2020-11- 9 | P4 | 2 | 820 | 103 | ||
| WASP-11 b | 2019-12- 6 | OUT1 | 4 | 820 | 103 | |
| 2019-12- 7 | IN1 | 5 | 820 | 101 | ||
| 2019-12- 7 | OUT1 | 5 | 820 | 106 | ||
| 2019-12- 8 | OUT1 | 4 | 820 | 102 | ||
| 2019-12-18 | IN2 | 5 | 756 | 109 | ||
| 2019-12-20 | OUT2 | 3 | 820 | 101 | ||
| 2019-12-21 | OUT2 | 1 | 820 | 110 | ||
| 2019-12-21 | OUT2 | 3 | 820 | 114 | ||
| 2019-12-24 | P4 | 2 | 916 | 96 | ||
| 2020- 7-25 | P4 | 2 | 820 | 129 | ||
| 2020-11- 8 | P4 | 2 | 820 | 86 | ||
| 2020-11-12 | P4 | 2 | 820 | 150 | ||
| WASP-52 b | 2019- 8-13 | IN1 | 4 | 820 | 93 | |
| 2019- 8-14 | OUT1 | 1 | 916 | 96 | ||
| 2019- 8-15 | OUT1 | 4 | 916 | 69 | ||
| 2019-10- 7 | OUT2 | 1 | 767 | 84 | ||
| 2019-10- 8 | OUT2 | 1 | 767 | 102 | ||
| 2019-10- 8 | IN2 | 5 | 639 | 76 | ||
| 2019-10-10 | OUT2 | 2 | 767 | 86 | ||
| 2019-10-10 | OUT2 | 3 | 767 | 95 | ||
| WASP-69 b | 2019- 9-20 | IN1 | 1 | 628 | 40 | |
| 2019-10-19 | OUT2 | 8 | 268 | 140 | ||
| 2019-10-20 | OUT2 | 4 | 532 | 273 | ||
| 2019-10-21 | IN2 | 7 | 437 | 215 | ||
| 2019-10-22 | OUT2 | 4 | 532 | 245 | ||
| WASP-76 b | 2020-10- 1 | OUT1 | 4 | 309 | 149 | |
| 2020-10- 2 | OUT1 | 8 | 320 | 124 | ||
| 2020-10- 2 | IN1 | 10 | 320 | 142 | ||
| 2020-10- 4 | IN2 | 10 | 320 | 123 | ||
| 2020-10- 5 | OUT2 | 4 | 309 | 166 | ||
| 2020-10- 6 | OUT2 | 4 | 309 | 131 | ||
| WASP-80 b | 2019- 5-17 | P4 | 10 | 309 | 109 | |
| 2019- 7-12 | OUT1 | 1 | 309 | 90 | ||
| 2019- 7-13 | OUT1 | 2 | 309 | 84 | ||
| 2019- 7-13 | IN1 | 10 | 309 | 118 | ||
| 2020- 7-23 | P4 | 4 | 511 | 130 | ||
| 2020- 8- 1 | P4 | 2 | 511 | 140 | ||
| WASP-127 b | 2020- 2-13 | OUT1 | 10 | 415 | 124 | |
| 2020- 2-14 | OUT1 | 5 | 415 | 141 | ||
| 2020- 2-15 | IN1 | 10 | 415 | 139 | ||
| 2020- 2-16 | OUT1 | 5 | 415 | 94 | ||
| 2020- 2-17 | OUT1 | 5 | 415 | 85 | ||
| 2020- 3- 6 | P4 | 2 | 415 | 47 | ||
| 2020-12- 8 | P4 | 2 | 511 | 195 | ||
| WASP-156 b | 2020-10-23 | OUT1 | 4 | 618 | 75 | |
| 2020-10-24 | IN1 | 6 | 554 | 63 | ||
| 2020-10-25 | OUT1 | 4 | 618 | 93 | ||
| 2020-11-20 | IN2 | 6 | 554 | 98 | ||
| 2020-11-20 | OUT2 | 4 | 554 | 91 | ||
| 2020-11-21 | OUT2 | 4 | 618 | 117 | ||
| 2020-11-22 | OUT2 | 4 | 618 | 122 | ||
| WASP-177 b | 2020- 8-18 | OUT1 | 3 | 831 | 84 | |
| 2020- 8-19 | OUT1 | 4 | 831 | 98 | ||
| 2020- 8-20 | IN1 | 4 | 831 | 87 | ||
| 2020- 8-20 | OUT1 | 4 | 831 | 86 | ||
| 2020- 8-21 | OUT1&2 | 4 | 831 | 86 | ||
| 2020- 8-22 | OUT2 | 1 | 831 | 40 | ||
| 2020- 8-23 | OUT2 | 4 | 831 | 65 | ||
| 2020- 8-23 | IN2 | 4 | 831 | 75 | ||
| 2020- 8-24 | OUT2 | 3 | 831 | 84 | ||
| 2020- 8-25 | OUT2 | 4 | 831 | 87 |